# 00footnotetext: Supplementary website: https://sites.google.com/view/hetgps

### v3

@@ 3 $r_{31}$ $r_{32}$ $r_{33}$
Generally speaking, multi-agent reinforcement learning is a concurrent optimization process that faces non-stationarity as agents constantly adapt to one another’s evolving policies (Zhang et al., 2019). HetGPS extends this challenge as agents must also adapt to a changing environment. Consequently, formal convergence guarantees to a *global* optimum remain an open question with regards to HetGPS in particular, but also MARL algorithms in general. However, recent theoretical work in environment co-design has established conditions under which convergence of bilevel optimization processes similar to HetGPS to *local* optima can be guaranteed, such as requiring sufficient smoothness of the environment dynamics and policy updates (Gao et al., 2024).
Despite the theoretical complexities inherent in multi-agent learning and bilevel optimization, as shown in Sec. 5 and App. Q, HetGPS demonstrates strong empirical stability even under adversarial initializations. This stability is expected, as it mirrors the practical success observed in related co-design and automated curriculum learning literature (Dennis et al., 2020; Gao et al., 2024).
@@
Appendix R Limitations and Open Questions
We list a number of limitations, open questions, and possible extensions.

### v4

@@ 3 $r_{31}$ $r_{32}$ $r_{33}$
Generally speaking, multi-agent reinforcement learning is a concurrent optimization process that faces non-stationarity as agents constantly adapt to one another’s evolving policies (Zhang et al., 2019). HetGPS extends this challenge as agents must also adapt to a changing environment. Consequently, formal convergence guarantees to a global optimum remain an open question with regards to HetGPS in particular, but also MARL algorithms in general. However, recent theoretical work in environment co-design has established conditions under which convergence of bilevel optimization processes similar to HetGPS to local optima can be guaranteed, such as requiring sufficient smoothness of the environment dynamics and policy updates (Gao et al., 2024).
Despite the theoretical complexities inherent in multi-agent learning and bilevel optimization, as shown in Sec. 5 and App. Q, HetGPS demonstrates strong empirical stability even under adversarial initializations. This stability is expected, as it mirrors the practical success observed in related co-design and automated curriculum learning literature (Dennis et al., 2020; Gao et al., 2024). However, we emphasize that empirical stability is not a guarantee of convergence. Although we did not identify such cases ourselves, it is possible that in some scenarios, HetGPS will oscillate rather than converge.
@@
Appendix R Best Practices, Limitations, and Open Questions
We list a number of limitations, open questions, and possible extensions. We then discuss best practices: where and how should our theoretical results be applied? What parameters should HetGPS optimize?
@@
R.4 Scope of the Curvature Analysis
The theoretical framework presented in Section 2 provides a precise characterization of the heterogeneity gain based on the curvature of reward aggregators. We provide an extended discussion of when we expect our theoretical predictions can, and cannot, be applied for deciding whether to use heterogeneous or homogeneous agent policies.
Symmetry and Monotonicity.
Our analysis hinges on the definition of generalized aggregators as symmetric and coordinate-wise non-decreasing. These assumptions are appropriate for studying emergent behavioral specialization among capability-identical agents. Symmetry ensures that agents (and tasks) are interchangeable ex-ante, isolating how the reward structure drives specialization. If symmetry is violated (e.g., due to inherently heterogeneous agent capabilities), heterogeneity is often trivially necessary. Monotonicity ensures a rational cooperative setting where increased effort does not decrease the reward.
Effort Constraints.
We define the feasible effort space over the closed unit simplex (where efforts sum $\leq 1$). However, because both the inner aggregators $T_{j}$ and the outer aggregator $U$ are non-decreasing, any optimal allocation—whether homogeneous or heterogeneous—will necessarily saturate the budget constraint (efforts sum =1=1). Therefore, our analysis focuses on this efficient frontier without loss of generality.
Constant-Sum Task Score Constraints.
It is crucial to clarify that the assumption of constant-sum task scores ($\sum_{j}T_{j}(a_{j})=C$) is specific only to Theorem 3.3, enabling the use of majorization to analyze the outer aggregator $U$. Theorems 3.1 and 3.2, and our sum-form aggregator analysis (App. F), do not rely on this.
When this assumption is violated, the analysis of $\Delta R$ involves a trade-off between the distribution of scores (influenced by curvature) and their total magnitude. Despite this complexity, our empirical findings (Section 5) suggest that the curvature analysis remains a robust heuristic for predicting the heterogeneity gain even when the task scores are variable-sum.
Reward functions that do not precisely follow the theory.
Our football experiments show that even when only part of the reward function adheres to our curvature theory (e.g., it is a sum $R(A)=R_{1}(A)+R_{2}(A)$ where $R_{1}$ is concave-convex and $R_{2}$ is a function with unclear curvature), our theoretical results may still predict the heterogeneity gain. We make no formal claims about the robustness of our predictions in such scenarios, but it is valuable to keep in mind that even if the entire reward function does not perfectly follow the theory, it may still be worthwhile to see what the concave-convex curvature test says about the part of it that does.